Hi
cleo wanted to buy 16 similarly priced shirts but was she was 8 dollars short.if she had bought 7 such shirts instead, she would have 19 dollars left.
what was the cost of 2 such shirts.
how much money did she have at first.
Thanks

Answer by richwmiller(17219)

16s=x+8
16s-8=x
7s=x-19
7s+19=x
16s-8=x
16s-8=7s+19
9s=27
s=3 each shirt cost $3
7*3=x-19
21=x-19
40=x cleo had $40
check
16s=x+8
16*3=40+8
48=48
7*3=40-19
21=21
ok